# A relative version of Bass’ theorem about finite-dimensional algebras

### v2

@@ Introduction
The main result about cotorsion pairs is the Eklof–Trlifaj theorem [6, Theorems 2 and 10], [10, Theorem 6.11 and Corollary 6.14]. The Eklof–Trlifaj theorem concerns the notion of a complete cotorsion pair, and its simplest formulation claims that every cotorsion pair generated by a set of modules is complete. Proving completeness of the flat cotorsion pair in the category of modules over an arbitrary ring was a major open problem, which became known under the name of the Flat Cover Conjecture. Posed by Enochs in [7] in 1981 (see also [18, Section 1.3]), it was proved by Bican–El Bashir–Enochs in [3] in 2001. One of the two proofs of the Flat Cover Conjecture presented in [3] was based on the Eklof–Trlifaj theorem.
@@
Specifically, the flaprojective cotorsion pair $(\mathsf{F},\mathsf{C})$ is formed by the following two classes of left $A$-modules. The right-hand class $\mathsf{C}$ consists of all left $A$-modules that are cotorsion as $R$-modules, i. e., the $A$-modules whose underlying $R$-module is cotorsion. The left-hand class $\mathsf{F}$ consists of all the left $A$-modules $F$ such that $\operatorname{Ext}^{1}_{A}(F,C)=0$ for all left $A$-modules $C$ that are cotorsion over $R$. The $A$-modules $F$ satisfying the latter condition are called flaprojective relative to $R$ (or $A/R$-flaprojective) [14, Section 1.6].
@@ 1. Preliminaries on Cotorsion Pairs
An introductory discussion of cotorsion pairs can be found in the introduction to [13]. Many examples of approximation sequences (for abelian groups, i. e., modules over the ring of integers $A=\mathbb{Z}$) are discussed in [12, Section 12]. The main reference source on cotorsion pairs in module categories is the book [10].
@@ 2. Cotorsion Modules
Using the facts that all projective left $R$-modules are flat and the class of flat modules is closed under kernels of epimorphisms in $R{\operatorname{\mathsf{--Mod}}}$, one can easily show that $\operatorname{Ext}_{R}^{n}(F,C)=0$ for all flat left $R$-modules $F$, all cotorsion left $R$-modules $C$, and all integers $n\geq 1$ (cf. [10, Lemma 5.24] and [15, Lemma 7.1]). It follows that the class of cotorsion $R$-modules is closed under cokernels of monomorphisms in $R{\operatorname{\mathsf{--Mod}}}$; it is also obviously closed under extensions and infinite products. We denote the class of flat left $R$-modules by $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}\subset R{\operatorname{\mathsf{--Mod}}}$ and the class of cotorsion left $R$-modules by $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}}\subset R{\operatorname{\mathsf{--Mod}}}$.
@@
This is [2, Theorem 5.1(2)]. See also [2, Theorem 1.2(2) or Proposition 4.8(2)] for essentially equivalent formulations of this theorem. A further discussion in category-theoretic context can be found in [15, Corollary 9.2 and Proposition 9.4]. ∎
@@
This is [11, Corollary 1.6.4]. The point is that any $A$-module can be embedded into a cotorsion $A$-module (in fact, all injective $A$-modules are cotorsion by definition). Therefore, any $A$-module $M$ has a coresolution (right resolution) by cotorsion $A$-modules, i. e., there is an exact sequence of $A$-modules
@@
This is [11, Lemma 1.3.4(a)] or [14, Lemma 1.6(a)]. One can use the $\operatorname{Ext}^{1}$-adjunction of change of rings as in [5, Proposition VI.4.1.3] (see Lemma 3.4 below). Alternatively, the following elementary argument works. A left $R$-module $C$ is cotorsion if and only if the functor $\operatorname{Hom}_{R}({-},C)$ takes every short exact sequence of flat left $R$-modules to a short exact sequence of abelian groups. Given a short exact sequence of flat left $R$-modules $0\longrightarrow H\longrightarrow G\longrightarrow F\longrightarrow 0$, we have a short exact sequence of flat left $A$-modules $0\longrightarrow A\otimes_{R}H\longrightarrow A\otimes_{R}G\longrightarrow A\otimes_{R}F\longrightarrow 0$. By assumption, the $A$-module $D$ is cotorsion, so the functor $\operatorname{Hom}_{A}({-},D)$ takes the latter short exact sequence of $A$-modules to a short exact sequence of abelian groups. It remains to recall the adjunction isomorphism of abelian groups $\operatorname{Hom}_{A}(A\otimes_{R}M,\>D)\simeq\operatorname{Hom}_{R}(M,D)$, which holds for any left $R$-module $M$. The assertion of the lemma is also a special case of the next Lemma 2.4. ∎
@@
This is [11, Lemma 1.3.3(a)] or [14, Lemma 1.5(a)]. The point is that the tensor product $K\otimes_{S}G$ is a flat left $R$-module for any flat left $S$-module $G$. Based on this observation, one can use the $\operatorname{Ext}^{1}$-adjunction of the tensor product and Hom functors as in [5, formula (4) in Section XVI.4], [11, formula (1.1) in Section 1.2], or [14, Lemma 1.1]. Alternatively, an elementary argument can be found in [14, proof of Lemma 1.5]. ∎
@@ 3. Flaprojective Modules
Let $R$ be an associative ring. It is easy to prove that the pair of classes of flat left $R$-modules and cotorsion left $R$-modules $(\mathsf{F},\mathsf{C})$ == $(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$, $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}})$ is a cotorsion pair in $R{\operatorname{\mathsf{--Mod}}}$[18, Lemma 3.4.1], [14, Lemma 1.3]. Proving that this pair of classes, called the flat cotorsion pair or the Enochs cotorsion pair [8], [10, Definition 5.18(i)]), is a complete cotorsion pair, is harder. There are two proofs of this result in the paper [3], one of them based on the following observation of Enochs.
@@
Now let $R\longrightarrow A$ be a homomorphism of associative rings. A left $A$-module $F$ is said to be flaprojective relative to $R$ (or $A/R$-flaprojective) [14, Section 1.6] if $\operatorname{Ext}^{1}_{A}(F,C)=0$ for all left $A$-modules $C$ that are cotorsion as left $R$-modules (i. e., the underlying left $R$-module of $C$ is cotorsion).
Since the class of cotorsion $R$-modules is closed under cokernels of monomorphisms (see the beginning of Section 2), so is the class of all $A$-modules that are cotorsion over $R$. All cotorsion $A$-modules are cotorsion over $R$ by Lemma 2.3; in particular, all injective $A$-modules are cotorsion over $R$. Using these observations, one can show that $\operatorname{Ext}^{n}_{A}(F,C)=0$ for all $A/R$-flaprojective left $R$-modules $F$, all left $A$-modules $C$ that are cotorsion over $R$, and all integers $n\geq 1$ (cf. [10, Lemma 5.24] and [15, Lemma 7.1]).
@@
(1) Since all cotorsion $A$-modules are cotorsion over $R$, it follows that all $A/R$-flaprojective AA-modules are flat [14, Remark 1.29]. Obviously, all projective $A$-modules are $A/R$-flaprojective; so the class of $A/R$-flaprojective AA-modules is, generally speaking, intermediate between the classes of projective and flat $A$-modules.
(2) If $A=R$ and $R\longrightarrow A$ is the identity ring homomoprhism, then the class of $A/R$-flaprojective AA-modules coincides with the class of flat $R$-modules [14, Example 1.28(1)].
@@
This is a particular case of [5, Proposition VI.4.1.3]. See also [11, formula (1.1) in Section 1.2] or [14, Lemma 1.1]. ∎
@@
This is [13, Proposition 3.3]. Notice that the left $A$-module $A^{+}=\operatorname{Hom}_{\mathbb{Z}}(A,\mathbb{Q}/\mathbb{Z})$ is injective, hence pure-injective, hence cotorsion; so by Lemma 2.3,  $A^{+}$ is a cotorsion left $R$-module and [13, Proposition 3.3] is applicable.
@@
Let $R$ be a left perfect ring in the sense of [1, Theorem P], i. e., all flat left $R$-modules are projective. Let $R\longrightarrow A$ be a ring homomorphism making $A$ a finitely generated projective right $R$-module. Then it follows from Corollary 3.7 that all flat left $A$-modules are projective, too. So $A$ is a left perfect ring as well.
@@
A homomorphism of associative rings $R\longrightarrow A$ is said to be a ring epimorphism if it is an epimorphism in the category of associative rings, or equivalently, if the functor of restriction of scalars $A{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ is fully faithful. A ring homomorphism $R\longrightarrow A$ is an epimorphism if and only if the natural map of abelian groups $N\otimes_{R}M\longrightarrow N\otimes_{A}M$ is an isomorphism for every right $A$-module NN and left $A$-module $M$. A ring epimorphism $R\longrightarrow A$ is said to be left flat if $A$ is a flat left $R$-module. We suggest the book [17, Sections XI.1–3] and the paper [9, Section 4] as the reference sources on (left flat and arbitrary) ring epimorphisms.

### v3

@@ Introduction
The main result about cotorsion pairs is the Eklof–Trlifaj theorem [6, Theorems 2 and 10], [10, Theorem 6.11 and Corollary 6.14]. The Eklof–Trlifaj theorem concerns the notion of a complete cotorsion pair, and its simplest formulation claims that every cotorsion pair generated by a set of modules is complete. Proving completeness of the flat cotorsion pair in the category of modules over an arbitrary ring was a major open problem, which became known under the name of the Flat Cover Conjecture. Posed by Enochs in [7] in 1981 (see also [21, Section 1.3]), it was proved by Bican–El Bashir–Enochs in [3] in 2001. One of the two proofs of the Flat Cover Conjecture presented in [3] was based on the Eklof–Trlifaj theorem. Subsequently it was understood [11, Corollary 6.6], [18, Theorem 4.5] that a proof of the Eklof–Trlifaj theorem can be also obtained as an application of Quillen’s small object argument [17, Lemma II.3.3].
@@
Specifically, the flaprojective cotorsion pair $(\mathsf{F},\mathsf{C})$ is formed by the following two classes of left $A$-modules. The right-hand class $\mathsf{C}$ consists of all left $A$-modules that are cotorsion as $R$-modules, i. e., the $A$-modules whose underlying $R$-module is cotorsion. The left-hand class $\mathsf{F}$ consists of all the left $A$-modules $F$ such that $\operatorname{Ext}^{1}_{A}(F,C)=0$ for all left $A$-modules $C$ that are cotorsion over $R$. The $A$-modules $F$ satisfying the latter condition are called flaprojective relative to $R$ (or $A/R$-flaprojective) [15, Section 1.6].
@@ 1. Preliminaries on Cotorsion Pairs
An introductory discussion of cotorsion pairs can be found in the introduction to [14]. Many examples of approximation sequences (for abelian groups, i. e., modules over the ring of integers $A=\mathbb{Z}$) are discussed in [13, Section 12]. The main reference source on cotorsion pairs in module categories is the book [10].
@@ 2. Cotorsion Modules
Using the facts that all projective left $R$-modules are flat and the class of flat modules is closed under kernels of epimorphisms in $R{\operatorname{\mathsf{--Mod}}}$, one can easily show that $\operatorname{Ext}_{R}^{n}(F,C)=0$ for all flat left $R$-modules $F$, all cotorsion left $R$-modules $C$, and all integers $n\geq 1$ (cf. [10, Lemma 5.24] and [16, Lemma 7.1]). It follows that the class of cotorsion $R$-modules is closed under cokernels of monomorphisms in $R{\operatorname{\mathsf{--Mod}}}$; it is also obviously closed under extensions and infinite products. We denote the class of flat left $R$-modules by $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}\subset R{\operatorname{\mathsf{--Mod}}}$ and the class of cotorsion left $R$-modules by $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}}\subset R{\operatorname{\mathsf{--Mod}}}$.
@@
This is [2, Theorem 5.1(2)]. See also [2, Theorem 1.2(2) or Proposition 4.8(2)] for essentially equivalent formulations of this theorem. A further discussion in category-theoretic context can be found in [16, Corollary 9.2 and Proposition 9.4]. ∎
@@
This is [12, Corollary 1.6.4]. The point is that any $A$-module can be embedded into a cotorsion $A$-module (in fact, all injective $A$-modules are cotorsion by definition). Therefore, any $A$-module $M$ has a coresolution (right resolution) by cotorsion $A$-modules, i. e., there is an exact sequence of $A$-modules
@@
This is [12, Lemma 1.3.4(a)] or [15, Lemma 1.6(a)]. One can use the $\operatorname{Ext}^{1}$-adjunction of change of rings as in [5, Proposition VI.4.1.3] (see Lemma 3.4 below). Alternatively, the following elementary argument works. A left $R$-module $C$ is cotorsion if and only if the functor $\operatorname{Hom}_{R}({-},C)$ takes every short exact sequence of flat left $R$-modules to a short exact sequence of abelian groups. Given a short exact sequence of flat left $R$-modules $0\longrightarrow H\longrightarrow G\longrightarrow F\longrightarrow 0$, we have a short exact sequence of flat left $A$-modules $0\longrightarrow A\otimes_{R}H\longrightarrow A\otimes_{R}G\longrightarrow A\otimes_{R}F\longrightarrow 0$. By assumption, the $A$-module $D$ is cotorsion, so the functor $\operatorname{Hom}_{A}({-},D)$ takes the latter short exact sequence of $A$-modules to a short exact sequence of abelian groups. It remains to recall the adjunction isomorphism of abelian groups $\operatorname{Hom}_{A}(A\otimes_{R}M,\>D)\simeq\operatorname{Hom}_{R}(M,D)$, which holds for any left $R$-module $M$. The assertion of the lemma is also a special case of the next Lemma 2.4. ∎
@@
This is [12, Lemma 1.3.3(a)]. The point is that the tensor product $K\otimes_{S}G$ is a flat left $R$-module for any flat left $S$-module $G$. Based on this observation, one can use the $\operatorname{Ext}^{1}$-adjunction of the tensor product and Hom functors as in [5, formula (4) in Section XVI.4], [12, formula (1.1) in Section 1.2], or [15, Lemma 1.1]. Alternatively, an elementary argument can be found in [15, proof of Lemma 1.5(a)]. ∎
@@ 3. Flaprojective Modules
Let $R$ be an associative ring. It is easy to prove that the pair of classes of flat left $R$-modules and cotorsion left $R$-modules $(\mathsf{F},\mathsf{C})$ == $(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$, $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}})$ is a cotorsion pair in $R{\operatorname{\mathsf{--Mod}}}$[21, Lemma 3.4.1], [15, Lemma 1.3]. Proving that this pair of classes, called the flat cotorsion pair or the Enochs cotorsion pair [8], [10, Definition 5.18(i)]), is a complete cotorsion pair, is harder. There are two proofs of this result in the paper [3], one of them based on the following observation of Enochs.
@@
Now let $R\longrightarrow A$ be a homomorphism of associative rings. A left $A$-module $F$ is said to be flaprojective relative to $R$ (or $A/R$-flaprojective) [15, Section 1.6] if $\operatorname{Ext}^{1}_{A}(F,C)=0$ for all left $A$-modules $C$ that are cotorsion as left $R$-modules (i. e., the underlying left $R$-module of $C$ is cotorsion).
Since the class of cotorsion $R$-modules is closed under cokernels of monomorphisms (see the beginning of Section 2), so is the class of all $A$-modules that are cotorsion over $R$. All cotorsion $A$-modules are cotorsion over $R$ by Lemma 2.3; in particular, all injective $A$-modules are cotorsion over $R$. Using these observations, one can show that $\operatorname{Ext}^{n}_{A}(F,C)=0$ for all $A/R$-flaprojective left $R$-modules $F$, all left $A$-modules $C$ that are cotorsion over $R$, and all integers $n\geq 1$ (cf. [10, Lemma 5.24] and [16, Lemma 7.1]).
@@
(1) Since all cotorsion $A$-modules are cotorsion over $R$, it follows that all $A/R$-flaprojective AA-modules are flat [15, Remark 1.29]. Obviously, all projective $A$-modules are $A/R$-flaprojective; so the class of $A/R$-flaprojective AA-modules is, generally speaking, intermediate between the classes of projective and flat $A$-modules.
(2) If $A=R$ and $R\longrightarrow A$ is the identity ring homomorphism, then the class of $A/R$-flaprojective AA-modules coincides with the class of flat $R$-modules [15, Example 1.28(1)].
@@
(4) The converse assertion to (3) is not true in general. For example, if $R=\mathbb{Z}$ is the ring of integers and $A=\mathbb{Q}$ is the field of rational numbers, then all $A$-modules are projective, hence they are also $A/R$-flaprojective by (1). But the ring $R$ is not perfect.
We are not aware of any characterization of the ring homomorphisms $R\longrightarrow A$ for which the class of $A/R$-flaprojective AA-modules coincides with the class of projective $A$-modules.
@@
This is a particular case of [5, Proposition VI.4.1.3]. See also [12, formula (1.1) in Section 1.2] or [15, Lemma 1.1]. ∎
@@
This is [14, Proposition 3.3]. Notice that the left $A$-module $A^{+}=\operatorname{Hom}_{\mathbb{Z}}(A,\mathbb{Q}/\mathbb{Z})$ is injective, hence pure-injective, hence cotorsion; so by Lemma 2.3,  $A^{+}$ is a cotorsion left $R$-module and [14, Proposition 3.3] is applicable.
@@
Let $R$ be a left perfect ring in the sense of [1, Theorem P]. Let $R\longrightarrow A$ be a ring homomorphism making $A$ a finitely generated projective right $R$-module. Then it follows from Corollary 3.7 that all flat left $A$-modules are projective. So $A$ is a left perfect ring as well.
@@
A homomorphism of associative rings $R\longrightarrow A$ is said to be a ring epimorphism if it is an epimorphism in the category of associative rings, or equivalently, if the functor of restriction of scalars $A{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ is fully faithful. A ring homomorphism $R\longrightarrow A$ is an epimorphism if and only if the natural map of abelian groups $N\otimes_{R}M\longrightarrow N\otimes_{A}M$ is an isomorphism for every right $A$-module NN and left $A$-module $M$. A ring epimorphism $R\longrightarrow A$ is said to be left flat if $A$ is a flat left $R$-module. We suggest the book [20, Sections XI.1–3] and the paper [9, Section 4] as the reference sources on (left flat and arbitrary) ring epimorphisms.
@@
Alternatively, one can simply say that for, for any flat $A$-module $F$, one has $F\simeq A\otimes_{R}F$ and $F$ is flat $R$-module; so the $A$-module $F$ is $A/R$-flaprojective by Theorem 3.5(b).